OpenStudy (anonymous):

Here's a matrix problem: "Suppose A is idempotent. Let C=1/√2 * <|> (note: each <> identify a column in the matrix). Show that C is an involution.

OpenStudy (anonymous):

If C is an involution C*C=I ¿right?

OpenStudy (anonymous):

yep !

OpenStudy (anonymous):

but what does <<I,A>|<A,-I>> mean? because my guess isi if you square it and use the fact that \[A^2=A\] it should pop right out

OpenStudy (anonymous):

Then, why don't follow this? C*C=1/√2 * <<I,A>| * 1/√2 * <<I,A>| =I, given that A is Idempotent, A*A=A

OpenStudy (anonymous):

The equallity should be fulfilled after perform all matrix operations

OpenStudy (anonymous):

@satellite73 <<I,A>|<A,-I>> means